# TiCoSS: Tightening the Coupling between Semantic Segmentation and Stereo Matching within A Joint Learning Framework

### v1

@@ Abstract
Semantic segmentation and stereo matching, respectively analogous to the ventral and dorsal streams in our human brain, are two key components of autonomous driving perception systems. Addressing these two tasks with separate networks is no longer the mainstream direction in developing computer vision algorithms, particularly with the recent advances in large vision models and embodied artificial intelligence. The trend is shifting towards combining them within a joint learning framework, especially emphasizing feature sharing between the two tasks. The major contributions of this study lie in comprehensively tightening the coupling between semantic segmentation and stereo matching. Specifically, this study introduces three novelties: (1) a tightly coupled, gated feature fusion strategy, (2) a hierarchical deep supervision strategy, and (3) a coupling tightening loss function. The combined use of these technical contributions results in TiCoSS, a state-of-the-art joint learning framework that simultaneously tackles semantic segmentation and stereo matching. Through extensive experiments on the KITTI and vKITTI2 datasets, along with qualitative and quantitative analyses, we validate the effectiveness of our developed strategies and loss function, and demonstrate its superior performance compared to prior arts, with a notable increase in mIoU by over 9%. Our source code will be publicly available at mias.group/TiCoSS upon publication.
@@ I-A Background
Therefore, in recent years, the simultaneous learning, deployment, and inference of both tasks have become a mainstream [20, 21, 16, 22, 23], because a unified joint learning framework can process the contextual and geometric information more comprehensively. It also enables end-to-end training of the entire system, capable of tackling the challenges posed by both tasks [16]. Consequently, this can lead to an improved overall performance in semantic segmentation and stereo matching, when compared to the models trained separately for each task [23]. Moreover, a unified framework can potentially reduce computational complexity and memory usage, as learning representations can likely be shared between both tasks [24]. This can be advantageous particularly for resource-constrained applications that require real-time performance.
@@ I-B Existing Challenges and Motivation
The performance of a feature-fusion semantic segmentation network largely depends on the manner in which heterogeneous feature fusion is conducted [25, 26]. A current bottleneck lies in the simplistic and indiscriminate fusion of heterogeneous features, which often causes conflicting learning representations and erroneous segmentation results [27]. Taking our previous work S3M-Net [16] as an example, its adopted feature fusion strategy essentially performs an element-wise summation between contextual and geometric feature maps at each stage. These feature maps are then directly fed into subsequent layers without excluding irrelevant information, leading to the loose coupling between the semantic segmentation and stereo matching tasks within the encoder. Furthermore, as the network goes deeper, such an indiscriminate feature fusion strategy tends to diminish the proportion of useful geometric features in the decoder’s input, potentially leading to unsatisfactory semantic segmentation performance.
Additionally, existing joint learning frameworks usually converge slowly during training, due to the vanishing gradient problem. A common solution is to employ deep supervision (DS) strategies that incorporate additional pathways to facilitate gradient propagation. Nonetheless, the existing DS strategies utilized in feature-fusion semantic segmentation networks typically overlook the potential interactions among auxiliary classifiers. This will result in limited improvement in semantic segmentation within our developed multi-task joint learning framework.
Finally, in the loss function aspect, previous joint learning frameworks, such as SegStereo[21] and SSNet [28], typically compute the losses for different tasks separately and supervise the entire training process by simply minimizing the weighted sum of these losses. This training function overlooks the complementarity between these two tasks at the output level, also resulting in their loose coupling. Regrettably, tightening the coupling between these two tasks, especially in the encoder and loss function aspects, has received relatively little attention in this research area and needs further investigation.
@@ I-C Contributions
Therefore, in this article, we introduce Tightly-Coupled Semantic Segmentation and Stereo Matching Network (TiCoSS), an end-to-end joint learning approach that focuses primarily on improving the coupling between stereo matching and semantic segmentation, which has not been emphasized in previous related studies. Building upon our previous work S3M-Net, our proposed TiCoSS introduces three new techniques: (1) a Tightly-Coupled, Gated Feature Fusion (TGF) strategy, which utilizes a series of Selective Inheritance Gates (SIGs) to propagate useful contextual and geometric information from the preceding layer to the current layer; (2) a Hierarchical Deep Supervision (HDS) strategy that uses the fused features with the highest resolution to guide deep supervision throughout each branch, as these features contain the most abundant local spatial details; (3) a novel Coupling Tightening (CT) loss, consisting of a widely used stereo matching loss presented in [9], the Semantic Consistency-Guided (SCG) loss in [16], a Disparity Inconsistency-Aware (DIA) loss that utilizes disparity estimations to refine segmentation results, and a Deep Supervision Consistency Constraint (DSCC) loss which employs the Kullback-Leibler (KL) divergence to improve prediction consistency across outputs from all deep supervision branches. These contributions collectively advance S3M-Net, resulting in TiCoSS, a new, powerful, and tightly-coupled joint learning framework that enables robust and accurate simultaneous semantic segmentation and stereo matching. Extensive experiments conducted on the vKITTI2 [29] and KITTI 2015 [30] datasets unequivocally demonstrate the effectiveness of the aforementioned contributions and the superior performance of TiCoSS over other state-of-the-art (SoTA) approaches.
@@ II-C Simultaneous Semantic Segmentation and Stereo Matching
Existing joint learning frameworks that simultaneously address these two tasks mainly focus on improving disparity accuracy by leveraging semantic information [21, 20, 12, 22, 23, 10]. However, discussions on utilizing disparity information to enhance semantic segmentation at the feature level for joint learning remain limited, except for the aforementioned “RGB-X semantic segmentation”. These prior arts often require extensive annotated training data or involve complex training strategies for joint learning. For example, SegStereo [21] and DispSegNet [20] necessitate an initial unsupervised training phase on the large-scale Cityscapes [48] dataset, followed by a subsequent supervised fine-tuning on the smaller KITTI [49, 30] datasets. Similarly, the approaches introduced in [12, 22, 23] require pre-training their spatial branches for stereo matching on the large-scale SceneFlow [50] dataset before fine-tuning both semantic and spatial branches on the KITTI [49, 30] datasets. DSNet [10] adopts a different joint learning strategy by alternating the training of the semantic segmentation and stereo matching networks, with each network being frozen during the training of the other. Nevertheless, achieving simultaneous convergence of the two networks can be challenging, as shared features are not learned in an end-to-end fashion. Recently, SSNet [28] uses a single encoder to extract shareable features for both tasks. However, as demonstrated in [51], such features may not be suitable for both dense prediction and geometric vision tasks. S3M-Net uses two separate encoding branches to achieve these two tasks simultaneously, but the weak coupling between these branches limits the integration of contextual and geometric information. Additionally, the source code (in PyTorch or TensorFlow) for these previous approaches is not publicly available, and re-implementing them carries the risk of introducing errors. In contrast to the aforementioned approaches, our proposed TiCoSS uses a tightly-coupled joint learning framework that effectively leverages both contextual and geometric information. Moreover, TiCoSS is trained in an end-to-end fashion and is capable of learning semantic segmentation and stereo matching simultaneously, even with limited training data.
@@ III-C Hierarchical Deep Supervision
After tightening the coupling between semantic segmentation and stereo matching during the feature encoding stage, we turn our focus towards the feature decoding aspect. We first revisit the deep supervision strategies employed in SNE-RoadSeg+ [39] and UNet 3+ [53]. The former applies deep supervision to the decoded features with the highest resolution, whereas the latter achieves this on the deepest decoded features at each resolution. Despite the effectiveness of these two prior approaches in addressing challenges such as vanishing gradients and slow model convergence, the deep supervision strategies employed by them are not comprehensive (the former emphasizes enhancing local details, while the latter focuses on improving multi-scale feature consistency). Ideally, they should be used in conjunction to complement each other for improved results. Therefore, our proposed HDS strategy combines the strengths of both methods and demonstrates superior performance compared to each individually.
@@
A straightforward way to combine the strengths of these two methods is to apply deep supervision simultaneously to both the main and side branches, enabling the network to leverage features from shallow layers (containing rich local details) and deep layers (being semantically strong). However, this strategy does not fully exploit the potential complementarity between the main and side auxiliary classifiers, since the decoded features are exclusively derived from adjacently-connected ones. To improve their interactions, we utilize the decoded, fused feature maps $\mathbf{F}_{1}^{F}$ (containing rich, fine-grained local spatial details that are essential for semantic segmentation) at the highest resolution to guide the feature decoding process in the side branches. Specifically, for the $l$-th auxiliary classifier within the side branch, we utilize a Feature Dynamic Alignment (FDA) block, which is composed of $l$ downsampling units to progressively align channel dimensions and spatial resolutions between a pair of features at different layers. Each downsampling unit comprises a $3\times 3$ convolutional layer with a stride of 2, followed by a batch normalization (BN) layer and a rectified linear unit (ReLU) activation layer. Compared to the simple max-pooling operation which may disrupt the original feature representation, our FDA achieves a smoother feature alignment. The feature maps obtained by downsampling $\mathbf{F}_{1}^{F}$ are then concatenated with the deepest decoded features at the corresponding layers. This downsampling output not only guides the decoding process but also serves as the input for the subsequent downsampling unit, thereby preserving fine-grained local details to the greatest extent possible. Since the outputs of the side branches are obtained by directly upsampling the deepest features at each layer, this strategy does not significantly impact training efficiency and memory usage. Moreover, the auxiliary classifiers of the main and side branches collaboratively provide additional pathways for gradients to propagate more efficiently from the output layers to their corresponding layers, thereby accelerating the convergence of our model.
@@ III-D1 Disparity Inconsistency-Aware Loss
In our previous work [16], the developed SCG loss focuses mainly on regions with semantic inconsistencies caused by occlusion and reflection. Given that relying solely on semantic consistency cannot sufficiently tighten the coupling between these two tasks at the output level, we define a weight matrix $\mathbf{W}\in\mathbb{R}^{H\times W}$ drawing from the concept of Left-Right Consistency in stereo matching [11], where
@@ IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 4090 GPU with a batch size of 1. We set the maximum disparity search range to 192 pixels. All images are cropped to 512 $\times$ 256 pixels before being processed by the network. We utilize the AdamW [55] optimizer for model training, with epsilon and weight decay set to $10^{-5}$ and $10^{-8}$, respectively. The initial learning rate is set to $2\times 10^{-4}$. Training lasts for 100,000 iterations on the vKITTI2 dataset and 20,000 iterations on the KITTI 2015 dataset. Standard data augmentation techniques are applied to improve model robustness.
@@ IV-C Evaluation Metrics
TABLE I: Quantitative comparisons of SoTA semantic segmentation networks on the KITTI 2015 [30] dataset. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance.
@@
TiCoSS (Ours) / Feature-Fusion 91.95 72.28 63.91 86.82 92.35 94.24 92.90
@@
TiCoSS (Ours) / Feature-Fusion 98.69 91.66 88.46 97.58 98.55 98.67 98.57
TABLE III: Comparisons of SoTA stereo matching network on the KITTI 2015 [30] and vKITTI2 [29] datasets. The symbol $\downarrow$ indicates that a lower value corresponds to better performance. The best results are shown in bold font.
Networks Publications KITTI 2015 [30] vKITTI2 [29]
@@
PSMNet [8] CVPR’18 0.74 16.12 2.61 0.68 10.31 3.77
LEA-Stereo [44] NeurIPS’20 0.83 18.67 3.22 0.83 13.33 4.84
RAFT-Stereo [9] 3DV’21 0.60 10.78 1.96 0.40 5.88 2.67
CRE-Stereo [45] CVPR’22 0.92 19.68 3.35 0.63 10.35 3.90
ACVNet [58] CVPR’22 0.68 13.93 2.10 0.61 9.41 3.45
PCW-Net [59] ECCV’22 0.70 14.81 2.43 0.63 9.45 3.49
IGEV-Stereo [60] CVPR’23 0.62 12.15 1.99 0.47 7.15 3.09
S3M-Net TIV’24 0.55 10.02 1.62 0.38 5.56 2.55
TiCoSS(Ours) / 0.53 10.09 1.58 0.36 5.49 2.58

### v2

@@ Abstract
Semantic segmentation and stereo matching, respectively analogous to the ventral and dorsal streams in our human brain, are two key components of autonomous driving perception systems. Addressing these two tasks with separate networks is no longer the mainstream direction in developing computer vision algorithms, particularly with the recent advances in large vision models and embodied artificial intelligence. The trend is shifting towards combining them within a joint learning framework, especially emphasizing feature sharing between the two tasks. The major contributions of this study lie in comprehensively tightening the coupling between semantic segmentation and stereo matching. Specifically, this study introduces three key novelties: (1) a tightly coupled, gated feature fusion strategy, (2) a hierarchical deep supervision strategy, and (3) a coupling tightening loss function. All these technical contributions help tightly couple semantic segmentation and stereo matching tasks, collectively resulting in TiCoSS, a state-of-the-art joint learning framework. Through extensive experiments on the KITTI and vKITTI2 datasets, along with both qualitative and quantitative analyses, we validate the effectiveness of our developed strategies and loss function. Our approach demonstrates superior performance compared to prior arts, with a notable increse in mIoU by over 9%. Our source code will be publicly available at mias.group/TiCoSS upon publication.
@@ I-A Background
Therefore, in recent years, the simultaneous learning, deployment, and inference of both tasks have become a mainstream [20, 21, 16, 22, 23], because a unified joint learning framework can process contextual and geometric information more comprehensively. It also enables end-to-end training of the entire system, capable of tackling the challenges posed by both tasks [16]. Consequently, this can enhance the overall performance in both semantic segmentation and stereo matching tasks, outperforming models trained separately for each task [23]. Moreover, a unified framework can potentially reduce computational complexity and memory usage, as learning representations can likely be shared between both tasks [24]. This can be advantageous particularly for resource-constrained applications that require real-time performance.
@@ I-B Existing Challenges and Motivation
The performance of a feature-fusion semantic segmentation network is heavily influenced by the employed strategy for heterogeneous feature fusion [25, 26]. Currently, the bottleneck lies in the simplistic and indiscriminate fusion of heterogeneous features, which often causes conflicting learning representations and erroneous segmentation results [27]. Taking our previous work S3M-Net [16] as an example, its adopted feature fusion strategy essentially performs an element-wise summation between contextual and geometric feature maps at each stage. These feature maps are then directly fed into subsequent layers without excluding irrelevant information, leading to a loose coupling between the semantic segmentation and stereo matching tasks within the encoder. Furthermore, as the network goes deeper, such an indiscriminate feature fusion strategy tends to diminish the proportion of useful geometric features in the decoder’s input [28], potentially leading to unsatisfactory semantic segmentation performance.
Additionally, due to the vanishing gradient problem [29], existing joint learning frameworks usually exhibit slow convergence during training. A common solution is to employ the deep supervision (DS) strategies that incorporate additional pathways to achieve gradient propagation [30]. Nonetheless, existing DS strategies employed in feature-fusion semantic segmentation networks typically overlook the potential interactions between the main and side auxiliary classifiers, failing to exploit the complementarity among all classifiers and can therefore restrict the improvements in semantic segmentation performance within our multi-task joint learning framework.
In the loss function aspect, previous joint learning frameworks, such as SegStereo[21] and SSNet [31], typically compute the losses for two tasks separately and supervise the entire training process by simply minimizing the weighted sum of these losses. Such loss function fails to exploit the complementarity between these two tasks at the output level, also resulting in their loose coupling.
Prior joint learning arts, such as S3M-Net [16] and DSNet [10], primarily focus on introducing a joint learning framework that performs semantic segmentation and stereo matching simultaneously. However, exploring the potential complementarity and tightening the coupling between these two tasks, particularly in the encoder, decoder and loss function aspects, have received relatively limited attention in this research area and calls for further investigation.
@@ I-C Contributions
To address the aforementioned limitations, we introduce Tightly-Coupled Semantic Segmentation and Stereo Matching Network (TiCoSS), an end-to-end joint learning approach that focuses primarily on improving the coupling between stereo matching and semantic segmentation tasks, which has not been emphasized in previous related studies. Our proposed TiCoSS introduces three new techniques: (1) a tightly-coupled, gated feature fusion (TGF) strategy, which utilizes a series of selective inheritance gates (SIGs) to propagate useful contextual and geometric information from the preceding layer to the current layer; (2) a hierarchical deep supervision (HDS) strategy that uses the highest resolution fused feature maps to guide deep supervision throughout each branch, as these features contain the most abundant local spatial details; (3) a novel coupling tightening (CT) loss, consisting of a widely used stereo matching loss presented in [9], the semantic consistency-guided (SCG) loss in [16], a disparity inconsistency-aware (DIA) loss that utilizes disparity estimations to refine segmentation results, and a deep supervision consistency constraint (DSCC) loss which employs the Kullback-Leibler (KL) divergence to improve prediction consistency across outputs from all deep supervision branches. These contributions collectively advance S3M-Net, resulting in TiCoSS, a new, powerful, and tightly-coupled joint learning framework that simultaneously performs robust and accurate semantic segmentation and stereo matching tasks. Extensive experiments conducted on the vKITTI2 [32] and KITTI 2015 [33] datasets unequivocally demonstrate the effectiveness of the aforementioned contributions and the superior performance of TiCoSS over other state-of-the-art (SoTA) approaches.
@@ II-C Simultaneous Semantic Segmentation and Stereo Matching
Existing joint learning frameworks that simultaneously address these two tasks mainly focus on improving disparity accuracy by leveraging semantic information [21, 20, 12, 22, 23, 10]. However, discussions on utilizing disparity information to enhance semantic segmentation at the feature level for joint learning remain limited, except for the aforementioned “RGB-X semantic segmentation”. These prior arts often require extensive annotated training data or involve complex training strategies for joint learning. For example, SegStereo [21] necessitates an initial unsupervised training phase on the large-scale Cityscapes [50] dataset, followed by a subsequent supervised fine-tuning on the smaller KITTI [51, 33] datasets. Similarly, the approaches introduced in [12, 22, 23] require pre-training their spatial branches for stereo matching on the large-scale SceneFlow [52] dataset before fine-tuning both semantic and spatial branches on the KITTI [51, 33] datasets. DSNet [10] adopts a different joint learning strategy by alternating the training of the semantic segmentation and stereo matching networks, with parameters of each network being frozen during the training of the other. Nevertheless, leveraging both contextual and geometric information can be challenging, as the shared features are not learned in an end-to-end manner. DispSegNet [20] utilizes an embedding learned from the semantic segmentation branch to refine the initial disparity estimation. RTS2Net [22] relies on coarse-to-fine estimations in a multi-stage fashion for accurate disparity estimation. However, both methods only achieve limited improvement in segmentation results since they fail to leverage the informative spatial details to enhance semantic segmentation. Recently, SSNet [31] uses a single encoder to extract shareable features for both tasks. However, as demonstrated in [53], such shareable features may not be suitable for both dense prediction and geometric vision tasks. S3M-Net uses two separate encoding branches to achieve these two tasks simultaneously, but the weak coupling between these branches limits the integration of contextual and geometric information. Additionally, the source code (in PyTorch or TensorFlow) for these previous approaches is not publicly available, and re-implementing them carries the risk of introducing errors. In contrast to the aforementioned approaches, our proposed TiCoSS uses a tightly-coupled joint learning framework that effectively leverages both contextual and geometric information. Moreover, TiCoSS is trained in an end-to-end manner and is capable of learning accurate and robust semantic segmentation and stereo matching tasks simultaneously, even with limited training data.
@@ III-C Hierarchical Deep Supervision
After tightening the coupling between semantic segmentation and stereo matching during the feature encoding stage, we turn our focus towards the feature decoding process. We first revisit the deep supervision strategies employed in SNE-RoadSeg+ [42] and UNet 3+ [55]. The former applies deep supervision to the decoded features with the highest resolution, whereas the latter achieves this on the deepest decoded features at each resolution. Despite the effectiveness of these two prior approaches in addressing challenges such as vanishing gradients and slow model convergence, the deep supervision strategies employed by them are not comprehensive (the former emphasizes enhancing local details, while the latter focuses on improving multi-scale feature consistency). Ideally, they should be used in conjunction to complement each other for improved results. Therefore, our proposed HDS strategy combines the strengths of both methods and demonstrates superior performance compared to each individually.
@@
A straightforward way to combine the strengths of these two methods is to apply deep supervision strategies simultaneously to both the main and side branches, enabling the network to leverage features from shallow layers (containing rich local details) and deep layers (being semantically strong). However, this method can not fully exploit the potential complementarity between the main and side auxiliary classifiers, since the decoded features are exclusively derived from adjacently-connected ones. Additionally, as discussed in Sect. III-B, decoded features in deep layers also lack informative spatial details which are typically contained in shallow layers, limiting the performance of DS strategies in our multi-task framework. Thus, to improve the interactions among auxiliary classifiers and provide local spatial details for decoded features in deep layers, we utilize the decoded, fused feature maps $\mathbf{F}_{1}^{F}$ (containing rich, fine-grained local spatial details that are essential for semantic segmentation) at the highest resolution to guide the feature decoding process in the side branches. Specifically, for the $l$-th auxiliary classifier within the side branch, we utilize a Feature Dynamic Alignment (FDA) block, which is composed of $l$ downsampling units to progressively align channel dimensions and spatial resolutions between a pair of features at different layers. Each downsampling unit comprises a $3\times 3$ convolutional layer with a stride of 2, followed by a batch normalization (BN) layer and a rectified linear unit (ReLU) activation layer. Compared to the simple max-pooling operation which may disrupt the original feature representation, our FDA achieves a smoother feature alignment. The feature maps obtained by downsampling $\mathbf{F}_{1}^{F}$ are then concatenated with the deepest decoded features at the corresponding layers. This downsampling output not only guides the decoding process but also serves as the input for the subsequent downsampling unit, thereby preserving fine-grained local details to the greatest extent possible. Since the outputs of the side branches are obtained by directly upsampling the deepest features at each layer, this strategy does not significantly impact training efficiency and memory usage. Moreover, the auxiliary classifiers of the main and side branches collaboratively provide additional pathways for gradients to propagate more efficiently from the output layers to their corresponding layers, thereby accelerating the convergence of our model.
@@ III-D1 Disparity Inconsistency-Aware Loss
In our previous study [16], the developed SCG loss focuses mainly on enforcing semantic consistency to reduce segmentation errors caused by occlusions, neglecting the fact that occlusions can also lead to inconsistent disparity estimations [56], which are unfortunately under-explored when designing the loss. Thus, we further incorporate the disparity inconsistency in it to form a more advanced loss function to train our TiCoSS. Specifically, we define a weight matrix $\mathbf{W}\in\mathbb{R}^{H\times W}$ drawing from the concept of Left-Right Consistency in stereo matching [11], where
@@ IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 4090 GPU with a batch size of 1. We set the maximum disparity search range to 192 pixels. All images are cropped to 512 $\times$ 256 pixels before being processed by the network. We utilize the AdamW [58] optimizer for model training, with epsilon and weight decay set to $10^{-8}$ and $10^{-5}$, respectively. The initial learning rate is set to $2\times 10^{-4}$. Training lasts for 100,000 iterations on the vKITTI2 dataset and 20,000 iterations on the KITTI 2015 dataset. Standard data augmentation techniques are applied to improve model robustness.
@@ IV-C Evaluation Metrics
TABLE I: Quantitative comparisons of SoTA semantic segmentation networks on the KITTI 2015 [33] dataset. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. Though more recent joint learning frameworks exist, the source code (in PyTorch or TensorFlow) for these previous approaches is not publicly available.
@@
TiCoSS (Ours) / 91.90 71.97 63.63 86.68 92.43 94.10 92.90
@@
TiCoSS (Ours) / 98.69 91.66 88.46 97.58 98.55 98.67 98.57
TABLE III: Quantitative comparisons of semantic segmentation networks on the KITTI 2015 [33] dataset, evaluated using the mmsegmentation framework [61] . The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance.
Methods mIoU (%) $\uparrow$ mAcc (%) $\uparrow$
S3M-Net [16] 36.01 41.81
RoadFormer [14] 41.55 49.09
TiCoSS (Ours) 44.91 52.99
TABLE IV: Comparisons of SoTA stereo matching network on the KITTI 2015 [33] and vKITTI2 [32] datasets. The symbol $\downarrow$ indicates that a lower value corresponds to better performance. The best results are shown in bold font.
Networks Publications vKITTI2 [32] KITTI 2015 [33]
@@
PSMNet [8] CVPR’18 0.68 10.31 3.77 0.74 16.12 2.61
LEA-Stereo [46] NeurIPS’20 0.83 13.33 4.84 0.83 18.67 3.22
RAFT-Stereo [9] 3DV’21 0.40 5.88 2.67 0.60 10.78 1.96
CRE-Stereo [47] CVPR’22 0.63 10.35 3.90 0.92 19.68 3.35
ACVNet [62] CVPR’22 0.61 9.41 3.45 0.68 13.93 2.10
PCW-Net [63] ECCV’22 0.63 9.45 3.49 0.70 14.81 2.43
IGEV-Stereo [64] CVPR’23 0.47 7.15 3.09 0.62 12.15 1.99
DispSegNet [20] RA-L’18 0.50 6.37 2.91 0.81 14.47 2.69
DSNet[10] ICRA’19 0.64 8.42 3.82 0.73 15.10 2.92
S3M-Net [16] TIV’24 0.38 5.56 2.55 0.55 10.02 1.62
TiCoSS(Ours) / 0.34 5.43 2.58 0.54 10.39 1.60